OpenStudy (anonymous):

Estimate the sum. 9/10+7/8 A) about 1 B) about 2 C)about 1/2 D) about 0

OpenStudy (anonymous):

you just did one like this, dont rely on the site for answers, how are we going to help you during your exam

OpenStudy (anonymous):

estimation = to the closest one - round up or down

OpenStudy (anonymous):

I'm terrible at math. I can't multiply at all nor can I divide

OpenStudy (anonymous):

And it isn't an exam.. its questions I got on my paper I have here

hero (hero):

Hint: Both terms are close to 1

hero (hero):

9/10 is close to 10/10 7/8 is close to 8/8

OpenStudy (anonymous):

there is no 8/8s on it a,b,c and d are what I have to choose

hero (hero):

But 10/10 = 1 8/8 = 1 1 + 1 = ?

OpenStudy (anonymous):

thank you I get it now >.<;